# Sparsity-Promoting Reachability Analysis and Optimization of Constrained Zonotopes

### v2

@@ VI-C Moving Horizon Estimation and Set-Valued State Estimation
Simulated states, MHE estimates, and feasible sets for the numerical example are given in Fig. 6, with Algorithm 5 used to compute the MHE estimates and feasible sets. The left sub-figure depicts the position results while the right sub-figure depicts the velocity results. The RMS position error from the measurements is 0.505 m and the RMS velocity error is 0.200 m/s. The RMS position error from the MHE solution is 0.224 m and the RMS velocity error is 0.066 m/s. Using the computer described in Fig. 5, the average and maximum solution times for this numerical example are $0.026$ ms and $0.057$ ms, respectively. The true state and MHE estimate always lie within the feasible set while the measurement may lie outside the feasible set. In this example, the widths of the feasible sets are generally much larger than the MHE RMS errors. By combining SVSE and MHE as has been done here, one can compute an optimal state estimate in the sense of (31) while simultaneously computing the feasible set $\mathcal{X}_{0}$ that contains all possible states of the system.
@@
Figure 6: Simulated position and velocity estimates and feasible sets using MHE.
@@ VI-D Safety Verification
Consider the safety verification problem described in Sec. V-C using a disturbed discrete-time double integrator system with $A$ and $B$ defined by (35). The unsafe set $\mathcal{O}$ only contains position states, so safety verification at time step $k$ entails verifying the infeasibility of $\mathcal{Z}=\mathcal{X}_{k}\cap_{\begin{bmatrix}I&0\end{bmatrix}}\mathcal{O}$. In ZonoOpt, this feasibility check is implemented by executing Algorithm 1 with $\tilde{P}=I$ and $\mathbf{q}=\mathbf{0}$.
@@
The results for this numerical example are shown in Fig. 7. Algorithm 1 finds a certificate of infeasibility for $\mathcal{X}_{k}\cap_{\begin{bmatrix}I&0\end{bmatrix}}\mathcal{O}$ for all time steps $k$, indicating that trajectories of the system are guaranteed to be safe. The sets $\mathcal{X}_{k}$ are color coded to show how many ADMM iterations were required to generate the infeasibility certificate. In the majority of cases, only a single iteration is required. Using the desktop computer described in Fig. 5, the average and maximum times to produce the infeasibility certificates are $0.0061$ ms and $0.013$ ms, respectively.
@@
Figure 7: Safety verification for the disturbed double integrator system with LQR state feedback control.
@@ VII Experimental Results
where $\mathcal{Z}_{\circ}$ is an outer approximation of a unit circle, taken here as an eight-sided zonotope. The velocity measurement error set $\mathcal{V}_{v}$ was computed analogously to $\mathcal{W}_{u}$, with the body frame velocity error set given as $\mathcal{V}^{b}_{v}\equiv[\mathbf{e}_{v}^{b}]$ where $\mathbf{e}_{v}^{b}]=\begin{bmatrix}[-e_{v},e_{v}]&[-\epsilon,\epsilon]\end{bmatrix}^{T}$. In this case, $e_{v}$ is the maximum error in the velocity derived from the wheel motor encoder measurements, and $0<\epsilon\ll 1$.
For the MHE, covariance matrices $Q$ and $R$ were estimated by transforming body frame covariance matrices into the lab frame as required for velocity and acceleration. Standard deviations are taken to be one half of the uncertainty bound, e.g., the body frame forward velocity standard deviation is $\sigma_{v}=\frac{1}{2}e_{v}$. The covariance matrices do not account for heading angle uncertainty contributions. As in Sec. VI-C, the initial inverse covariance matrix was taken to be $P^{-1}_{N}=0$.
@@
TABLE II: Parameters for SVSE/MHE experimental implementation.
@@
Parameters for the experimental implementation are given in Table II. The heading uncertainty $e_{\theta}$ accounts for both initialization error and errors in the IMU’s heading computation. The velocity uncertainty $e_{v}$ is dominated by errors due to asynchronous sampling between the IMU and wheel motor encoder measurements. Acceleration uncertainty bounds were estimated from inspection of recorded acceleration data.
Results for this experiment are displayed in Fig. 8. Here, the MHE position estimates $\hat{\mathbf{x}}_{p,k}$ and projections of the SVSE bounding sets $\mathcal{X}_{k}$ onto the position states are overlaid on top of overhead camera images of the laboratory space. The motion capture positions are given by $\mathbf{x}_{p,k}$ and are taken to be ground truth. The position measurement feedback for the SVSE/MHE is given by $\mathbf{y}_{p,k}$. Because there is significant overlap between adjacent $\mathcal{X}_{k}$, only every tenth iteration is plotted. Effectiveness of the SVSE is verified by the observation that $\mathbf{x}_{p,k}\in[I\;0]\mathcal{X}_{k}\;\forall k$. The MHE significantly improved position estimation accuracy when compared to the position measurements. The RMS error for the position measurements $\mathbf{y}_{p,k}$ was 0.31 m, while the RMS error for the MHE position estimates $\hat{\mathbf{x}}_{p,k}$ was 0.11 m.

### v3

@@ VI-C Moving Horizon Estimation and Set-Valued State Estimation
Simulated states, MHE estimates, and feasible sets for the numerical example are given in Fig. 7, with Algorithm 3 used to compute the MHE estimates and feasible sets. The left sub-figure depicts the position results while the right sub-figure depicts the velocity results. The RMS position error from the measurements is 0.505 m and the RMS velocity error is 0.200 m/s. The RMS position error from the MHE solution is 0.224 m and the RMS velocity error is 0.066 m/s. Using the computer described in Fig. 5, the mean solution time for this example was $0.27$ ms (min: $0.05$ ms, max: $0.58$ ms). The true state and MHE estimate always lie within the feasible set while the measurement may lie outside the feasible set. In this example, the widths of the feasible sets are generally much larger than the MHE RMS errors. By combining SVSE and MHE as has been done here, one can compute an optimal state estimate in the sense of (33) while simultaneously computing the feasible set $\mathcal{X}_{0}$ that contains all possible states of the system.
@@
Figure 7: Simulated position and velocity estimates and feasible sets for a double integrator system using combined SVSE and MHE. Given an initial set containing the state of the system, the blue feasible sets are guaranteed to contain the system state. The MHE provides a best estimate of the system state within the blue feasible sets.
@@ VI-D Safety Verification
Consider the safety verification problem described in Sec. V-C using a disturbed discrete-time double integrator system with $A$ and $B$ defined by (37). This scenario is depicted in Fig. 8. The unsafe set $\mathcal{O}$ only contains position states, so safety verification at time step $k$ entails verifying the infeasibility of $\mathcal{Z}=\mathcal{X}_{k}\cap_{\begin{bmatrix}I&0\end{bmatrix}}\mathcal{O}$. In ZonoOpt, this feasibility check is implemented by executing Algorithm 1 with $\tilde{P}=I$ and $\mathbf{q}=\mathbf{0}$.
@@
Fig. 8 shows the results for this numerical example. Algorithm 1 finds a certificate of infeasibility for $\mathcal{X}_{k}\cap_{\begin{bmatrix}I&0\end{bmatrix}}\mathcal{O}$ for all time steps $k$, indicating that trajectories of the system are guaranteed to be safe. The sets $\mathcal{X}_{k}$ are color coded to show how many ADMM iterations were required to generate the infeasibility certificate. In the majority of cases, only a single iteration is required. Using the desktop computer described in Fig. 5, the mean time to produce an infeasibility certificate was $0.058$ ms (min: $0.024$ ms, max: $0.129$ ms).
@@
Figure 8: Safety verification for the disturbed double integrator system with LQR state feedback control. The $\mathcal{X}_{k}$ are the projections of the reachable sets of the system, $\mathcal{O}$ is the unsafe set, and the $\mathbf{x}^{r}_{k}$ are the state references. Thm. 1 is used to prove that there exists no trajectory of the system that intersects the unsafe set in a handful of ADMM iterations.
@@ VII-B Implementation Details
where $\mathcal{Z}_{\circ}$ is an outer approximation of a unit circle, taken here as an eight-sided zonotope, which over-approximates the area of the set by about $5\%$. The velocity measurement error set $\mathcal{V}_{v}$ was computed analogously to $\mathcal{W}_{u}$, with the body frame velocity error set given as $\mathcal{V}^{b}_{v}\equiv[\mathbf{e}_{v}^{b}]$ where $\mathbf{e}_{v}^{b}]=\begin{bmatrix}[-e_{v},e_{v}]&[-\epsilon,\epsilon]\end{bmatrix}^{T}$. In this case, $e_{v}$ is the maximum error in the velocity derived from the wheel motor encoder measurements, and $0<\epsilon\ll 1$.
For the MHE, covariance matrices $Q$ and $R$ were estimated by transforming body frame covariance matrices into the lab frame as required for velocity and acceleration. Standard deviations are taken to be one half of the uncertainty bound, e.g., the body frame forward velocity standard deviation is $\sigma_{v}=\frac{1}{2}e_{v}$. The covariance matrices do not account for heading angle uncertainty contributions. The initial inverse covariance matrix was taken to be $P^{-1}_{N}=0$.
@@
TABLE III: Parameters for SVSE/MHE experimental implementation.
@@
Parameters for the experimental implementation are given in Table III. The heading uncertainty $e_{\theta}$ accounts for both initialization error and errors in the IMU’s heading computation. The velocity uncertainty $e_{v}$ is dominated by errors due to asynchronous sampling between the IMU and wheel motor encoder measurements. Acceleration uncertainty bounds were estimated from inspection of recorded acceleration data.
[…]
Figure 10: Time series plot of motion capture positions $\mathbf{x}_{p,k}$, position measurements $\mathbf{y}_{p,k}$, MHE estimates $\hat{\mathbf{x}}_{p,k}$, and SVSE bounds $\mathcal{X}_{k}$ for the first 75 time steps of the combined MHE and SVSE experiment.
VII-C Results
Results for this experiment are displayed in Fig. 9. Here, the MHE position estimates $\hat{\mathbf{x}}_{p,k}$ and projections of the SVSE bounding sets $\mathcal{X}_{k}$ onto the position states are overlaid on top of overhead camera images of the laboratory space. The motion capture positions are given by $\mathbf{x}_{p,k}$ and are taken to be ground truth. The position measurement feedback for the SVSE/MHE is given by $\mathbf{y}_{p,k}$. Because there is significant overlap between adjacent $\mathcal{X}_{k}$, only every tenth iteration is plotted in Fig. 9. A time-series plot of the same data is given in Fig. 10. For readability, only the first 75 time steps (of 400) are displayed. Effectiveness of the SVSE is verified by the observation that $\mathbf{x}_{p,k}\in[I\;0]\mathcal{X}_{k}\;\forall k$. The MHE significantly improved position estimation accuracy when compared to the position measurements. The RMS error for the position measurements $\mathbf{y}_{p,k}$ was 0.31 m, while the RMS error for the MHE position estimates $\hat{\mathbf{x}}_{p,k}$ was 0.11 m.
Problem formulation and solution time statistics for the SVSE/MHE experiment are given in Tab. IV. The MHE converged quickly in this experiment. This is because the optimal estimate was generally in the interior of the SVSE feasible set, so $\bm{\xi}_{k}$ iterates (26a) were typically feasible and few ADMM iterations were needed to achieve dual convergence. The main computational burden in this experiment was computing the bounding box over-approximations, as these each required eight invocations of Alg. 1, and optimal solutions were on the boundary of the SVSE feasible set.
TABLE IV: Problem formulation and solution time statistics for SVSE/MHE experiment.
unit median min max
SVSE (Alg. 3) time ms 8.58 0.067 13.35
MHE (Alg. 1) time ms 2.06 0.20 3.71
MHE (Alg. 1) iterations counts 2 1 4
Bounding box time ms 49.99 31.54 69.69